# PixelVLA: Advancing Pixel-level Understanding in Vision-Language-Action Model

### v1

@@ Abstract
Vision-Language-Action models (VLAs) are emerging as powerful tools for learning generalizable visuomotor control policies. However, current VLAs are mostly trained on large-scale image–text–action data and remain limited in two key ways: (i) they struggle with pixel-level scene understanding, and (ii) they rely heavily on textual prompts, which reduces their flexibility in real-world settings. To address these challenges, we introduce PixelVLA, the first VLA model designed to support both pixel-level reasoning and multimodal prompting with text and visual inputs. Our approach is built on a new visuomotor instruction tuning framework that integrates a multiscale pixel-aware encoder with a visual prompting encoder. To train PixelVLA effectively, we further propose a two-stage automated annotation pipeline that generates Pixel-160K, a large-scale dataset with pixel-level annotations derived from existing robot data. Experiments on three standard VLA benchmarks and two VLA model variants show that PixelVLA improves manipulation success rates by $10.1\%\sim 28.7\%$ over OpenVLA, while requiring only $1.5\%$ of its pretraining cost. These results demonstrate that PixelVLA can be integrated into existing VLAs to enable more accurate, efficient, and versatile robot control in complex environments. The code will be released as open source.
@@ 1 Introduction
Traditional robotic policy learning methods (Brohan et al. (2022); Liang et al. (2024); Chi et al. (2023)) rely heavily on task-specific demonstration datasets (James et al. (2020); Liu et al. (2023a)), which limits their ability to generalize to out-of-distribution (OOD) tasks. In contrast, vision-language-action models (VLAs) (Brohan et al. (2023); Kim et al. (2024); Black et al. (2024)) leverage large-scale robot datasets together with pre-trained vision-language models (VLMs), achieving stronger generalization and instruction-following capabilities. For example, RT-2 (Brohan et al. (2023)) integrates internet-scale VLMs with robotic control, enabling semantic reasoning and manipulation of novel objects. Similarly, OpenVLA (Kim et al. (2024)) leverages Prismatic VLM (Karamcheti et al. (2024)) as backbone to conduct large-scale training on the OXE dataset (O’Neill et al. (2024)), leading to significant improvement in OOD generalization.
@@
Despite recent progress, as shown in Fig. 1 (a), most VLAs (Kim et al. (2024); Yang et al. (2025); Shi et al. (2025)) inherit from VLMs that process observations only at the image level, lacking fine-grained pixel-level understanding. This gap limits spatial reasoning and weakens OOD generalization. Pixel-level comprehension, in contrast, enables precise object perception and richer spatial awareness, key for robust manipulation in diverse environments. The second limitation lies in prompting. Most VLAs depend solely on textual instructions, which overlook subtle visual cues and constrain multimodal human–robot interaction (Jiang et al. (2023); Zheng et al. (2024)). To explore visual prompting in VLAs, TraceVLA (Zheng et al. (2024)) improves spatial-temporal awareness with visual traces, and LLaRA (Li et al. (2025a)) encodes object locations within textual prompts to enhance region-level understanding. Nevertheless, these approaches still face challenges in achieving fine-grained pixel-level understanding and effectively integrating diverse multimodal prompts (e.g., points, lines, regions, masks) (Jiang et al. (2023); Wu et al. (2024b); Dong et al. (2024)).
@@
Inspired by the successful visual instruction tuning in VLMs (Liu et al. (2023b); Karamcheti et al. (2024)), we introduce a novel visuomotor instruction tuning framework to train our VLA models. This framework is designed to significantly enhance the pixel-level understanding capabilities of VLAs and empower them to effectively process multimodal visuomotor control prompts. However, current robotic datasets (O’Neill et al. (2024); Khazatsky et al. (2024)) lack multimodal prompts and pixel-level annotations. Meanwhile, directly employing existing VLMs and open-set segmentation models (Karamcheti et al. (2024); Liu et al. (2024); Sun et al. (2024)) to extract visual prompts and pixel-level annotations proves to be ineffective. This is due to a significant domain gap between their pre-training data and robotic data, as well as the cluttered and low-quality nature of robotic images.
@@
To tackle the above challenges, as presented in Fig. 1 (b), we introduce PixelVLA in this paper, the first vision-language-action model that achieves both pixel-level understanding and multimodal prompting. The model architecture of PixelVLA comprises a pre-trained VLMs as backbone, a visual prompting encoder, a multiscale pixel-aware encoder and a continuous action decoder. Specifically, in PixelVLA, we introduce a lightweight visual prompting encoder to process the diverse visual prompts (e.g., points, lines, regions, masks). Subsequently, a novel multiscale pixel-aware encoder is designed to generate pixel-aware embeddings to inject pixel-level understanding into VLAs. Furthermore, we develop a continuous action representation decoder that leverages pixel-level understanding to capture fine-grained action details based on the hidden states of VLMs.
@@
  • We present PixelVLA, a novel vision-language-action model enabling pixel-level understanding while supporting both textual and visual prompts. In PixelVLA, we introduce a lightweight visual prompting encoder to process diverse visual prompts, a novel multiscale pixel-aware encoder for pixel-level understanding injection, and a continuous action decoder to generate robotic action.
  • We design a novel two-stage automated annotation pipeline to effectively create a pixel-level visuomotor instruction tuning dataset form the publicly available robot datasets, called Pixel-160K, where the pipeline comprises the gripper-aware region proposal stage and the multimodal object segmentation stage.
@@ 2 Related Work
Vision-Language-Action Models. Vision-language-action models (VLAs) (Team et al. (2025); Brohan et al. (2024); Black et al. (2024); Ding et al. (2024); Fan et al. (2025)) have propelled robotic manipulation forward by endowing robots with the ability to understand and execute language-based instructions in diverse visual environments. Trained on numerous robot episodes, OpenVLA (Kim et al. (2024)) enables zero-shot control and adaptation for various robots. Building on the foundational capabilities of OpenVLA, various approaches have been proposed to advance robotic manipulation, such as SpatialVLA (Qu et al. (2025)) and ECoT (Zawalski et al. (2024)). Most prior VLAs focus on innovations in visual processing for robotic manipulation, such as integrating lightweight vision-language models with diffusion policy decoders for precise actions (Wen et al. (2025)) and introducing visual chain-of-thought (CoT) reasoning mechanisms for visual planning (Zhao et al. (2025)). Nevertheless, they primarily process visual information at the image level, lacking the ability to perform detailed pixel-level visual processing required for precise robotic manipulation.
Visual Instruction Tuning in VLAs. Visual Instruction Tuning (Zhu et al. (2023); Liu et al. (2023b); Rasheed et al. (2024)) is generally divided into two steps, which are modality alignment and instruction optimization, respectively. This strategy also serves as the core paradigm for realizing multimodal capabilities in VLAs (Li et al. (2025a), Zheng et al. (2024); Zawalski et al. (2024)). For example, TraceVLA (Zheng et al. (2024)) introduces visual trace prompting to enhance spatial-temporal awareness in VLAs. In contrast, LLaRA (Li et al. (2025a)) reformulates the robot action policy as visuo-textual conversations through visuomotor instruction tuning and RoVI (Li et al. (2025b)) develops an object-centric visual instruction paradigm with symbolic sketches. However, to address various visuomotor control challenges, adapting visual instruction tuning for VLAs remains a major constraint.
@@ 3 Problem Definition: Visuomotor Instruction Tuning
Inspired by the effectiveness of visual instruction tuning in VLMs (Liu et al. (2023b); Rasheed et al. (2024); Wu et al. (2024a)), we aim to adapt a similar process for VLAs to tackle diverse visuomotor control challenges (e.g., various multimodal prompts) and achieve pixel-level understanding. Similar to LLaRA (Li et al. (2025a)), we formalize this paradigm as Visuomotor Instruction Tuning. Specifically, following OpenVLA (Kim et al. (2024)), given a series of image observations $\mathbf{X}=\{\mathbf{x}^{t}\in\mathbb{R}^{H\times W\times 3}\}^{T}_{t=1}$ and a language instruction $\mathbf{L}$, the VLA model $\mathcal{F}_{\theta}(\cdot)$ can generate a series of robotic actions $\mathbf{A}=\{\mathbf{a}^{t}\in\mathbb{R}^{7}\}^{T}_{t=1}$, i.e., $\mathbf{a}^{t}=\mathcal{F}_{\theta}(\mathbf{x}^{t},\mathbf{L})$. For an episode of length $T$, the likelihood of successfully completing the task through an action sequence $\mathbf{A}$ can be calculated as:
@@
Figure 2: Overview of the PixelVLA architecture. The model integrates three novel components: (1) a visual prompting encoder for processing input diverse visual prompts; (2) a multiscale pixel-aware encoder that injects pixel-level information into token embeddings; and (3) a continuous action decoder to predict 7D robot actions. PixelVLA enhances fine-grained pixel-level spatial understanding and multimodal prompt responsiveness, enabling more precise manipulation policies in visually complex scenarios.
@@ 4 The Proposed Method
As illustrated in Fig. 2, we present the architecture of the proposed PixelVLA to achieve pixel-level understanding and accommodate both textual and visual prompts. Specifically, PixelVLA integrates a novel multiscale pixel-aware encoder (Sec. 4.1) that infuses pixel-level understanding into VLAs through tokenized representations, a visual prompting encoder for handling diverse visual prompts (Sec. 4.1), and a continuous action decoder (Sec. 4.1) for accurate robotic action prediction. In addition, a automated annotation generation pipeline and a pixel-annotated visuomotor instruction tuning dataset Pixel-160K are presented in Sec. 4.2. Subsequently, we introduce the proposed visuomotor instruction tuning procedure for training PixelVLA in Sec. 4.3.
@@ 4.1 PixelVLA Architecture
Current VLAs (Black et al. (2024); Kim et al. (2024); Wen et al. (2025)) are typically pre-trained on large-scale image-instruction-action robotic datasets (O’Neill et al. (2024); Wu et al. (2024c)). Architecturally built upon VLMs, these models process single or multi-view images along with textual instructions. However, this foundation inherently restricts their ability to achieve pixel-level understanding or respond to detailed visual prompts, resulting in constraining VLAs for spatial comprehension and object perception.
To address these architectural constraints, as illustrated in Fig. 2, we present a novel VLA model, namely PixelVLA. Specifically, PixelVLA integrates four main parts: (1) a vision encoder and MLP projector for visual embedding extraction, (2) a visual prompting encoder and a multiscale pixel-aware encoder for accommodating visual prompts and pixel-level understanding injection, (3) a LLM backbone and (4) a continuous action decoder for non-discrete robot action prediction. Following OpenVLA (Kim et al. (2024)), we preliminarily build our PixelVLA on Prismatic-7B VLM (Karamcheti et al. (2024)), where a Llama 2-7B (Touvron et al. (2023)) is employed as LLM backbone. Additionally, the vision encoder of PixelVLA consists of pre-trained DinoV2 (Oquab et al. (2023)) and SigLIP (Zhai et al. (2023)) models, and a lightweight 2-layer MLP projector is utilized to map the output features of the vision encoder into the input space of LLM. Notably, to accommodate visual prompts, we incorporate a lightweight prompt encoder from SAM (Kirillov et al. (2023)) as the visual prompting encoder within PixelVLA.
Multiscale Pixel-aware Encoder. To extract pixel-level information from multiscale image features and encode the spatial positional information of visual prompts, we propose a multiscale pixel-aware encoder designed to generate both pixel-aware embeddings and prompt-aware embeddings. Specifically, given the initial image observation $\mathbf{x}^{0}\in\mathbb{R}^{H\times W\times 3}$, PixelVLA first extracts the multi-level visual features $\mathbf{F}^{0}_{v}=\{\mathbf{f}_{v}^{0,i}\in\mathbb{R}^{H_{i}\times W_{i}\times D_{i}}\}_{i=1}^{L}$ through its vision encoder, where $L$ is the number of selected feature levels in the SigLIP vision encoder. As illustrated in Fig. 2 (a), the multiscale pixel-aware encoder leverages the features $\mathbf{F}^{0}_{v}$ and a pixel-aware mask input $\mathbf{p}^{0}\in\mathbb{R}^{H\times W}$ to compute the pixel-aware embeddings $\mathbf{E}_{p}^{0}\in\mathbb{R}^{N_{p}\times D}$. Here, $N_{p}$ is the length of pixel-aware embeddings and $D$ denotes the feature dimension of LLM. Specifically, the pixel-aware embeddings $\mathbf{E}^{0}_{p}$ can be computed as follows:
@@
where $\mathrm{MLP}(\cdot)$ is a multilayer perceptron (MLP) layer and $\Gamma^{i}(\cdot)$ denotes the linear projection in the $i$-th linear projector. To preserve the spatial positional information of the visual prompts, we first employ the visual prompting encoder to extract the visual prompting features and use a MLP layer to produce the prompt-aware embeddings $\mathbf{E}^{0}_{s}\in\mathbb{R}^{N_{s}\times D}$, where $N_{s}$ is the length of embeddings.
Continuous Action Decoder. Most existing VLA models (Kim et al. (2024); Zheng et al. (2024); Li et al. (2024a)) adapt autoregressive generation to predict sequential action tokens based on the pre-trained VLM backbone. In contrast, following $\pi_{0}$ (Black et al. (2024)), we develop a continuous action decoder that directly predicts continuous action representations, leveraging pixel-level understanding to capture fine-grained action details. Specifically, as illustrated in Fig. 2 (b), the hidden states $\mathbf{F}^{t}\in\mathbb{R}^{N_{s}\times D}$ from the last layer of LLM backbone are sequentially processed by a linear projector, $N_{r}$ ResNet blocks and a MLP projector to obtain the actions $\mathbf{A}\in\mathbb{R}^{N_{c}\times 7}$. Here, $N_{s}$ denotes the sequence length of the LLM backbone, while $N_{c}$ represents the chunk size used in action chunking (Zhao et al. (2023)). In this way, we can effectively preserves the pixel-level understanding learned by the pre-trained VLM backbone while enabling the continuous action decoder to incorporate these features directly into the next-token prediction.

### v2

@@ Abstract
Vision-Language-Action models (VLAs) are emerging as powerful tools for learning generalizable visuomotor control policies. However, current VLAs are mostly trained on large-scale image–text–action data and remain limited in two key ways: (i) they struggle with pixel-level scene understanding, and (ii) they rely heavily on textual prompts, which reduces their flexibility in real-world settings. To address these challenges, we introduce PixelVLA, the first VLA model designed to support both pixel-level reasoning and multimodal prompting with text and visual inputs. Our approach is built on a new visuomotor instruction tuning framework that integrates a multiscale pixel-aware encoder with a visual prompt-aware encoder. To train PixelVLA effectively, we further propose a two-stage automated annotation pipeline that generates Pixel-160K, a large-scale dataset with pixel-level annotations derived from existing robot data. Experiments on three standard VLA benchmarks and two VLA model variants show that PixelVLA improves manipulation success rates by $10.1\%\sim 28.7\%$ over OpenVLA, while requiring only $1.5\%$ of its pretraining cost. These results demonstrate that PixelVLA can be integrated into existing VLAs to enable more accurate, efficient, and versatile robot control in complex environments. Project page: https://wenqiliang.github.io/PixelVLA/.
@@ 1 Introduction
Traditional robotic policy learning methods (Brohan et al. (2022); Liang et al. (2024); Chi et al. (2023); Dong et al. (2026)) rely heavily on task-specific demonstration datasets (James et al. (2020); Liu et al. (2023a)), which limits their ability to generalize to out-of-distribution (OOD) tasks. In contrast, vision-language-action models (VLAs) (Brohan et al. (2023); Kim et al. (2024); Black et al. (2024)) leverage large-scale robot datasets together with pre-trained vision-language models (VLMs), achieving stronger generalization and instruction-following capabilities. For example, RT-2 (Brohan et al. (2023)) integrates internet-scale VLMs with robotic control, enabling semantic reasoning and manipulation of novel objects. Similarly, OpenVLA (Kim et al. (2024)) leverages Prismatic VLM (Karamcheti et al. (2024)) as backbone to conduct large-scale training on the OXE dataset (O’Neill et al. (2024)), leading to significant improvement in OOD generalization.
@@
Despite recent progress, as shown in Fig. 1 (a), most VLAs (Kim et al. (2024); Han et al. (2026); Shi et al. (2025)) inherit from VLMs that process observations only at the image level, lacking fine-grained pixel-level understanding. This gap limits spatial reasoning and weakens OOD generalization. In contrast, pixel-level comprehension has already been successfully validated in VLMs (Ren et al. (2024); Zhang et al. (2024b)) and enables precise object perception and richer spatial awareness, which are key for robust manipulation in diverse environments. The second limitation lies in prompting. Most VLAs depend solely on textual instructions, which overlook subtle visual cues and constrain spatial awareness (Ranasinghe et al. (2024)) and multimodal human–robot interaction (Jiang et al. (2023); Zheng et al. (2024); Li et al. (2026)). To explore visual prompting in VLAs, TraceVLA (Zheng et al. (2024)) improves spatial-temporal awareness with visual traces, and LLaRA (Li et al. (2025b)) encodes object locations within textual prompts to enhance region-level understanding. Nevertheless, these approaches still face challenges in achieving fine-grained pixel-level understanding and effectively integrating diverse multimodal prompts (e.g., points, lines, regions, masks) (Wu et al. (2024b); Dong et al. (2024)).
@@
Inspired by the successful visual instruction tuning in VLMs (Liu et al. (2023b); Karamcheti et al. (2024)), we introduce a novel visuomotor instruction tuning framework to train our VLA models. This framework is designed to significantly enhance the pixel-level understanding capabilities of VLAs and empower them to effectively process multimodal visuomotor control prompts. However, current robotic datasets (O’Neill et al. (2024); Khazatsky et al. (2024)) lack multimodal prompts and pixel-level annotations. Meanwhile, directly employing existing VLMs and open-set segmentation models (Karamcheti et al. (2024); Liu et al. (2024a)) to extract visual prompts and pixel-level annotations proves to be ineffective. This is due to a significant domain gap between their pre-training data and robotic data, as well as the cluttered and low-quality nature of robotic images.
@@
To tackle the above challenges, as presented in Fig. 1 (b), we introduce PixelVLA in this paper, the first vision-language-action model that achieves both pixel-level understanding and multimodal prompting. The model architecture of PixelVLA comprises a pre-trained VLMs as backbone, a visual prompt-aware encoder, a multiscale pixel-aware encoder and a continuous action decoder. Specifically, in PixelVLA, we introduce a lightweight visual prompt-aware encoder to process the diverse visual prompts (e.g., points, lines, regions, masks). Subsequently, a novel multiscale pixel-aware encoder is designed to generate pixel-aware embeddings to inject pixel-level understanding into VLAs. Furthermore, we develop a continuous action representation decoder that leverages pixel-level understanding to capture fine-grained action details based on the hidden states of VLMs.
@@
  • We present PixelVLA, a novel vision-language-action model enabling pixel-level understanding while supporting both textual and visual prompts. In PixelVLA, we introduce a lightweight visual prompt-aware encoder to process diverse visual prompts, a novel multiscale pixel-aware encoder for pixel-level understanding injection, and a continuous action decoder to generate robotic action.
  • We design a novel two-stage automated annotation pipeline to effectively create a pixel-level visuomotor instruction tuning dataset from the publicly available robot datasets, called Pixel-160K, where the pipeline comprises the gripper-aware region proposal stage and the multimodal object segmentation stage.
@@ 2 Related Work
Vision-Language-Action Models. Vision-language-action models (VLAs) (Team et al. (2025); Brohan et al. (2023); Black et al. (2024); Ding et al. (2024); Fan et al. (2025)) have propelled robotic manipulation forward by endowing robots with the ability to understand and execute language-based instructions in diverse visual environments. Trained on numerous robot episodes, OpenVLA (Kim et al. (2024)) enables zero-shot control and adaptation for various robots. Building on the foundational capabilities of OpenVLA, various approaches have been proposed to advance robotic manipulation, such as SpatialVLA (Qu et al. (2025)) and ECoT (Zawalski et al. (2024)). Most prior VLAs focus on innovations in visual processing for robotic manipulation, such as introducing visual chain-of-thought (CoT) reasoning mechanisms for visual planning (Zhao et al. (2025)). Nevertheless, they primarily process visual information at the image level, lacking the ability to perform detailed pixel-level visual processing required for precise robotic manipulation.
Visual Prompting in VLMs. Visual prompting methods (Zhang et al. (2023); Ren et al. (2024); Wu et al. (2024a); Zhang et al. (2024b)) have recently emerged as a complementary paradigm to textual prompting, allowing models to accept more fine-grained supervision in the form of region-level (Guo et al. (2024)) and even pixel-level instructions (Ma et al. (2024); Rasheed et al. (2024)) over multimodal inputs. Shikra (Chen et al. (2023)) extends MLLMs with a simple vision–encoder–LLM architecture that treats spatial coordinates as natural language tokens. Ferret (You et al. (2023)) enhances region-level grounding in MLLMs through a hybrid region representation and a spatial-aware visual sampler that supports diverse region inputs, while Ferret-v2 (Zhang et al. (2024a)) further introduces any-resolution grounding and multi-granularity visual encoding, leading to improved fine-grained visual understanding and localization over prior MLLMs. LocVLM (Ranasinghe et al. (2024)) proposes an image-space coordinate–based instruction fine-tuning framework for MLLMs that explores coordinate representations and spatial objectives to inject spatial awareness. However, despite these advances, robust pixel-level understanding in VLA frameworks remains challenging, especially when aligning fine-grained spatial cues with continuous, high-precision action control.
Visual Instruction Tuning in VLAs. Visual Instruction Tuning (Zhu et al. (2023); Liu et al. (2023b); Rasheed et al. (2024)) is generally divided into two steps, which are modality alignment and instruction optimization, respectively. This strategy also serves as the core paradigm for realizing multimodal capabilities in VLAs (Li et al. (2025b), Zheng et al. (2024); Zawalski et al. (2024)). For example, TraceVLA (Zheng et al. (2024)) introduces visual trace prompting to enhance spatial-temporal awareness in VLAs. In contrast, LLaRA (Li et al. (2025b)) reformulates the robot action policy as visuo-textual conversations through visuomotor instruction tuning and RoVI (Li et al. (2025c)) develops an object-centric visual instruction paradigm with symbolic sketches. However, to address various visuomotor control challenges, adapting visual instruction tuning for VLAs remains a major constraint.
@@ 3 Problem Definition: Visuomotor Instruction Tuning
Inspired by the effectiveness of visual instruction tuning in VLMs (Liu et al. (2023b); Rasheed et al. (2024); Wu et al. (2024a); Kang et al. (2025); Li et al. (2025a)), we aim to adapt a similar process for VLAs to tackle diverse visuomotor control challenges (e.g., various multimodal prompts) and achieve pixel-level understanding. Similar to LLaRA (Li et al. (2025b)), we formalize this paradigm as Visuomotor Instruction Tuning. Specifically, following OpenVLA (Kim et al. (2024)), given a series of image observations $\mathbf{X}=\{\mathbf{x}^{t}\in\mathbb{R}^{H\times W\times 3}\}^{T}_{t=1}$ and a language instruction $\mathbf{L}$, the VLA model $\mathcal{F}_{\theta}(\cdot)$ can generate a series of robotic actions $\mathbf{A}=\{\mathbf{a}^{t}\in\mathbb{R}^{7}\}^{T}_{t=1}$, i.e., $\mathbf{a}^{t}=\mathcal{F}_{\theta}(\mathbf{x}^{t},\mathbf{L})$. For an episode of length $T$, the likelihood of successfully completing the task through an action sequence $\mathbf{A}$ can be calculated as:
@@
Figure 2: Overview of the PixelVLA architecture. The model integrates three novel components: (1) a visual prompt-aware encoder for processing input diverse visual prompts; (2) a multiscale pixel-aware encoder that injects pixel-level information into token embeddings; and (3) a continuous action decoder to predict 7D robot actions. PixelVLA enhances fine-grained pixel-level spatial understanding and multimodal prompt responsiveness, enabling more precise manipulation policies in visually complex scenarios.
@@ 4 The Proposed Method
As illustrated in Fig. 2, we present the architecture of the proposed PixelVLA to achieve pixel-level understanding and accommodate both textual and visual prompts. Specifically, PixelVLA integrates a novel multiscale pixel-aware encoder (Sec. 4.1) that infuses pixel-level understanding into VLAs through tokenized representations, a visual prompt-aware encoder for handling diverse visual prompts (Sec. 4.1), and a continuous action decoder (Sec. 4.1) for accurate robotic action prediction. In addition, an automated annotation generation pipeline and a pixel-annotated visuomotor instruction tuning dataset Pixel-160K are presented in Sec. 4.2. Subsequently, we introduce the proposed visuomotor instruction tuning procedure for training PixelVLA in Sec. 4.3.
@@ 4.1 PixelVLA Architecture
Current VLAs (Black et al. (2024); Kim et al. (2024); Wang et al. (2026b; a)) are typically pre-trained on large-scale image-instruction-action robotic datasets (O’Neill et al. (2024); Wu et al. (2024c)). Architecturally built upon VLMs, these models process single or multi-view images along with textual instructions. However, this foundation inherently restricts their ability to achieve pixel-level understanding or respond to detailed visual prompts, resulting in constraining VLAs for spatial comprehension and object perception.
To address these architectural constraints, as illustrated in Fig. 2, we present a novel VLA model, namely PixelVLA. Specifically, PixelVLA integrates four main parts: (1) a vision encoder and MLP projector for visual embedding extraction, (2) a visual prompt-aware encoder and a multiscale pixel-aware encoder for accommodating visual prompts and pixel-level understanding injection, (3) a LLM backbone and (4) a continuous action decoder for non-discrete robot action prediction. Following OpenVLA (Kim et al. (2024)), we preliminarily build our PixelVLA on Prismatic-7B VLM (Karamcheti et al. (2024)), where a Llama 2-7B (Touvron et al. (2023)) is employed as LLM backbone. The vision encoder of PixelVLA consists of pre-trained DinoV2 (Oquab et al. (2023)) and SigLIP (Zhai et al. (2023)) models, and a lightweight 2-layer MLP projector is utilized to map the output features of the vision encoder into the input space of LLM.
Multiscale Pixel-aware Encoder. To extract pixel-level information from multiscale image features and encode the spatial positional information of visual prompts, we propose a multiscale pixel-aware encoder designed to generate both pixel-aware embeddings and prompt-aware embeddings. Specifically, as described in Eq. (2), for each training sample $\{\mathbf{x}^{0},\mathbf{p}^{0},\mathbf{L}^{0},\mathbf{V}^{0}\}$ drawn from Pixel-160K, PixelVLA first encodes the image observation $\mathbf{x}^{0}\in\mathbb{R}^{H\times W\times 3}$ with the SigLIP vision encoder to obtain multi-level visual features $\mathbf{F}^{0}_{v}=\{\mathbf{f}_{v}^{0,i}\in\mathbb{R}^{H_{i}\times W_{i}\times D_{i}}\}_{i=1}^{L}$, where $L$ denotes the number of selected feature levels. As illustrated in Fig. 2 (a), the multiscale pixel-aware encoder leverages the features $\mathbf{F}^{0}_{v}$ and a pixel-aware mask input $\mathbf{p}^{0}\in\mathbb{R}^{H\times W}$ to compute the pixel-aware embeddings $\mathbf{E}_{p}^{0}\in\mathbb{R}^{N_{p}\times D}$. Here, $N_{p}$ is the length of pixel-aware embeddings and $D$ denotes the feature dimension of LLM. Specifically, the pixel-aware embeddings $\mathbf{E}^{0}_{p}$ can be computed as follows:
@@
where $\mathrm{MLP}(\cdot)$ is a multilayer perceptron (MLP) layer and $\Gamma^{i}(\cdot)$ denotes the linear projection in the $i$-th linear projector. Supervised by the action prediction loss, PixelVLA learns to associate the pixel-level information encoded in these pixel-aware embeddings $\mathbf{E}^{0}_{p}$ with action generation, thereby enhancing the VLA backbone with pixel-level understanding.
Visual Prompt-aware Encoder. As shown in Fig. 2(b), we adopt a lightweight prompt encoder similar to that in SAM (Kirillov et al. (2023)) and integrate it into PixelVLA as the visual prompting encoder. Concretely, the user-provided prompts $\mathbf{V}^{0}\in\mathbb{R}^{H\times W}$ are first converted into continuous positional embeddings based on their normalized image coordinates, and then combined with learned prompt-type embeddings to produce prompt features $\mathbf{F}^{0}_{s}\in\mathbb{R}^{N_{s}\times D_{s}}$, where $N_{s}$ is the embedding length and $D_{s}$ is the feature dimension. These features $\mathbf{F}^{0}_{s}$ are further transformed by an MLP to obtain the final prompt-aware embeddings $\mathbf{E}^{0}_{s}\in\mathbb{R}^{N_{s}\times D}$. Since each embedding is explicitly tied to a specific location or region in the image via its coordinate-based positional embedding, the spatial positional information of the visual prompts is preserved throughout the encoding process.
Continuous Action Decoder. Most existing VLA models (Kim et al. (2024); Zheng et al. (2024); Li et al. (2024b)) adapt autoregressive generation to predict sequential action tokens based on the pre-trained VLM backbone. In contrast, following $\pi_{0}$ (Black et al. (2024)), we develop a continuous action decoder that directly predicts continuous action representations, leveraging pixel-level understanding to capture fine-grained action details. Specifically, as illustrated in Fig. 2 (b), the hidden states $\mathbf{F}^{t}\in\mathbb{R}^{N_{s}\times D}$ from the last layer of LLM backbone are sequentially processed by a linear projector, $N_{r}$ ResNet blocks and a MLP projector to obtain the actions $\mathbf{A}\in\mathbb{R}^{N_{c}\times 7}$. Here, $N_{s}$ denotes the sequence length of the LLM backbone, while $N_{c}$ represents the chunk size used in action chunking (Zhao et al. (2023)). In this way, we can effectively preserve the pixel-level understanding learned by the pre-trained VLM backbone while enabling the continuous action decoder to incorporate these features directly into the continuous action prediction. The resulting continuous actions are then used to compute an L1 regression loss that supervises the training process.